OpenStudy (anonymous):

use the quadratic formula to solve 4x^2+3x=8 round your answer to the nearest tenth

OpenStudy (anonymous):

please I really need help!

OpenStudy (anonymous):

do you know the quadraatic formula?

OpenStudy (anonymous):

yes its x=-b \sqrt{b^2-4ac} all over 2a. my teacher didn't teach me what to do if theres a negative number in the square root sign. I really need help. i also keep forgetting the steps.

OpenStudy (anonymous):

first thing is to write the quadratic equation as ax^2+bx+c=0,that is move the 8 to the left handside: 4x^2+3x-8=0 now a=4,b=3,c=-8 now compute b^2-4ac

OpenStudy (anonymous):

it shouldn't be negative

OpenStudy (anonymous):

yes it should be negative because you have to subtract the 8. 3+or- sqrt{9-119} all over 8

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

you haven't calculated it yet?whats the answer??

OpenStudy (anonymous):

sqrt{(3^2)-4(4)(-8)} = 11 something

OpenStudy (anonymous):

sqrt {9 +128}

OpenStudy (anonymous):

-3+or-sqrt{-119} all over 8. then i don't know what to do from here.

OpenStudy (anonymous):

therefore +/- 1.08808...

OpenStudy (anonymous):

inside your square root you have \[\sqrt{9-4(4)(-8)}\] the negative before the 4ac and the negative on the -8 cancel out leaving 9 + 128 not 9-128 which is what you have calculated

OpenStudy (anonymous):

oh ok I think i understand but can you still help me.

OpenStudy (anonymous):

@martinez the roots aren't opposites of eachother

OpenStudy (anonymous):

what now i am totally confused. please explain.

OpenStudy (anonymous):

i told you to calculate b^2-4ac,do it first

OpenStudy (anonymous):

-3+or-sqrt{-119} all over 8. then i don't know what to do from here. is that even right?

OpenStudy (anonymous):

no more -119 it is -3 + or - sqrt{137} all over 8. Since it is a quadratic with a + and - you will get 1.1 and -1.8

OpenStudy (anonymous):

ok how did you do that?

OpenStudy (anonymous):

\[-b \pm \sqrt{b ^{2}} -4(a)(c)\] all over 2a

OpenStudy (anonymous):

so \[(-3\pm \sqrt{3^{2}-4(4)(-8)})/2(4)\]

OpenStudy (anonymous):

(-3 + sqrt{137} ) / 8 && (-3 - sqrt{137} ) / 8

OpenStudy (anonymous):

ok i see that. thank you very much you just really helped me.

OpenStudy (anonymous):

also on your earlier question about what your teacher wants you to do for a negative number under the square root sign. At this moment nothing, but if you want to show off an answer (most likely a mistake) use i. i stands for a non real number \[\sqrt{-1}\] and -1 multiplied by anything negative becomes positive. Therefore adding i to any answer that has a negative under the square root makes the answer under the square root a positive times i.

OpenStudy (anonymous):

ok so where would i put the i and the out side of the square root sign or the inside?

OpenStudy (anonymous):

Anything I say from this point forward you should discredit as relevant to your homework until a trigonometry/ precalc/ calc class. That being said: the form would go \[a \pm bi\] where a and b are variables that have no relevance I just chose them as such and where this particular equation would be \[(-3 \pm i \sqrt{137}) / 8\]

OpenStudy (anonymous):

ok cool that helps a lot. thanks your wonderful.